A test has 20 single selection questions with five choices and each question worth 5 credits....

  1. A test has 20 single selection questions with five choices and each question worth 5 credits. In order to avoid student randomly guess the answer during the test and still get some credits, what will be reasonable cridits to be taken out for wrong answer from the student in each question? Prove it theoretically
Answer #1

P(getting correct answer while guessing) = 1 / Number of choices

= 1/5

When reasonable credit is taken out for wrong answer, the expected credit for guessing must be 0

Assume that k is the credit taken out for wrong answer

Expected value = 5 x 1/5 - k x 4/5 = 0

1 - 4k/5 = 0

k = 5/4

= 1.25 credits

Reasonable credits to be taken out for wrong answer from the student in each question = 1.25